6. [20] Let A, B e Cnxn such that A2 = A and B2 = B. Prove that if (A + B)2 = A + B, then AB is the zero matrix 0 Rnxn.
Answer #1

Julian : LA?= A), 162-B7 |(A+B)* = AB] expanding (A + B)² (A + B)² = 42+ B² + AB + BA (as in matrices A B & B 4, are differenIf you are satisfied with the solution, can upvote ? ?